OpenStudy (anonymous):

Graph each pair of parametric equations. 1.x = 3 sin3t y = 3 cos3t

ganeshie8 (ganeshie8):

square them and add

OpenStudy (anonymous):

x² = 9sin² 3t y² = 9cos² 3t

OpenStudy (anonymous):

x² + y² = 9sin² 3t + 9cos² 3t = 9(sin² 3t + cos² 3t)

ganeshie8 (ganeshie8):

yess, next use the identity : \(\sin^2 \theta + \cos^2 \theta = 1\)

OpenStudy (anonymous):

sin² 3t + cos² 3t = 1, so... x² + y² = 9

ganeshie8 (ganeshie8):

yes, does that look familiar to u ?

OpenStudy (anonymous):

yess

ganeshie8 (ganeshie8):

draw it :)

OpenStudy (anonymous):

so now i graph

ganeshie8 (ganeshie8):

yes, just draw it, and after that u can figure out the direction

OpenStudy (anonymous):

i think i need help is it graphed as a circle

ganeshie8 (ganeshie8):

you're right ! \(x^2+ y^2 = r^2\) is an equation of circle wid radius \(r\) and center at origin

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

wat about the direction ? when u graph a parametric curve, u must specify direction also :)

ganeshie8 (ganeshie8):

x = 3 sin3t y = 3 cos3t when t = 0, x = 0, y = 3 when t = pi/6, x = 3, y = 0

ganeshie8 (ganeshie8):